OpenStudy (anonymous):

can sombody solve this quadratic and explain it to me?? 3x^3-18X^2+24X=0

OpenStudy (anonymous):

try factoring out 3x from all the terms

OpenStudy (anonymous):

and then plugging into the quadratic formula??

OpenStudy (anonymous):

nah, you don't really need the quadratic formula for this

OpenStudy (anonymous):

\[3x^3 - 18x^2 + 24x = 0\\3x(x^2 - 6x + 8) = 0\\3x(x-4)(x-2) = 0\]

OpenStudy (anonymous):

3x^3-18X^2+24X=0 3x( x^2 - 6x +8) = 0

OpenStudy (anonymous):

then you can factor further then stuff in the parenthesis

OpenStudy (anonymous):

I guess I should have mentioned that your solving to find x is that possible??

OpenStudy (anonymous):

since it was x^3 in the equation, there should be three x values as solutions, just like x^2 equations have two solutions

OpenStudy (anonymous):

@mewi2671 solving to find x for that equation? Simple. Any of those factors could be 0 to cause the entire function to be 0. So, try for each one. \[3x=0\\x=0\] \[x-4=0\\x=4\] \[x-2=0\\x=2\]

OpenStudy (anonymous):

Thank you all very very much!!!!

OpenStudy (anonymous):

I like that approach... good stuff :)